# thanks: M.C. and D.P. contributed equally to this workthanks: M.C. and D.P. contributed equally to this work

### v2

@@ II Regular black holes
A final caveat is in order before discussing the RBH metrics we consider. The latter are all regular in the sense of having finite curvature invariants $R\equiv g^{\mu\nu}R_{\mu\nu}$, $R_{\mu\nu}R^{\mu\nu}$, and ${\cal K}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. However, a more stringent criterion for regularity is that of geodesic completeness, which does not necessarily imply finiteness of curvature invariants and viceversa. A number of “popular” RBHs have indeed been shown to have finite curvature invariants but to be geodesically incomplete Zhou and Modesto (2023). This includes the well-known Bardeen and Hayward RBHs, which are among the ones we shall consider here. Nevertheless, given the significant interest in these metrics, the fact that they are widely taken as prototypes for RBHs, and our phenomenological goal of going beyond Schwarzschild PBHs, we will take these space-times into consideration, while cautioning the reader about the above issues, and therefore that these metrics should be considered nothing more than phenomenological toy models at this stage. Note, in addition, that the stability of RBHs featuring inner horizons is currently a matter of debate in the literature Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023).
@@ II.1 Bardeen black hole
where, in terms of the BH mass $M$, the regularizing parameter satisfies $\ell\leq\sqrt{16/27}M\sim 0.77\,M$ in order for the metric to describe a BH and not a horizonless object. Note that the Schwarzschild metric function is recovered in the $\ell\to 0$ limit. A perhaps physically more motivated choice is to express quantities in units of the horizon radius $r_{H}$, defined as the largest root of the equation $f(r_{H})=0$, in which case the regularizing parameter is subject to the constraint $\ell\lesssim 0.70\,r_{H}$. In order to obtain this limit we have computed the solution to $f(r_{H})=0$ fixing $M=1$, in order to extrapolate $r_{H}(\ell)$, before analyzing for which real values of the parameter $n$ the equation $\ell=nr_{H}(\ell)$ admits solutions.
@@ II.2 Hayward black hole
If expressed in terms of BH mass $M$, the regularizing parameter for the Hayward BH is subject to the same limit as that of the Bardeen BH, i.e. $\ell\leq\sqrt{16/27}\,,M$. On the other hand, if expressed in terms of the more physically motivated horizon radius, the limit is instead $\ell\lesssim 0.57\,r_{H}$. We note that the Schwarzschild metric function is recovered in the $\ell\to 0$ limit.
@@ II.3 Culetu-Ghosh-Simpson-Visser black hole
II.4 Hayward-like black hole
For all of the three RBHs discussed earlier, the temperature decreases with increasing regularizing parameter, while converging to the Schwarzschild value $T_{\text{Sch}}=1/8\pi M$ as $\ell\to 0$, see Fig. 1. It is therefore not unreasonable to expect that evaporation constraints on $f_{\text{pbh}}$ may be weakened if DM consists of these PRBHs (this will of course be checked explicitly later). We have verified that such a behaviour of temperature versus regularizing parameter holds for most PRBHs of phenomenological interest considered in the literature.  [^3] This is actually not unrelated to the fact that most RBHs, and more generally hairy BH solutions, feature shadows whose size decreases relative to the size of the Schwarzschild BH shadow, i.e. $3\sqrt{3}M$, see Ref. Vagnozzi et al. (2023) for further more detailed discussions (see also Refs. Vertogradov and Övgün (2024a, b); Vertogradov et al. (2024)). In fact, the equations governing BH temperature and shadow size bear some resemblance, so an increase/decrease in one is expected to lead to an increase/decrease in the other. Ref. Vagnozzi et al. (2023) also presents a few space-times whose shadow radii increase with increasing hair parameter, and for these we would expect the temperatures to increase as well. However, for a few of these space-times we explicitly tested (which, we stress, are phenomenological in nature), the Teukolsky equation used to calculate the greybody factors turned out to be numerically very complicated, which is why we opted for proposing the equally phenomenological Hayward-like BH discussed here. For phenomenological reasons, especially in light of our expectation (verified a posteriori) that such a behaviour would lead to weaker bounds on $f_{\text{pbh}}$ and thereby the asteroid mass window further opening, it could therefore be desirable to also contemplate a case where the BH temperature actually increases with increasing regularizing parameter.
With the previous considerations in mind, here we introduce a new RBH solution metric which somewhat resembles the Hayward BH. This Hayward-like RBH is characterized by the following metric function:
$f_{\text{H-l}}(r)=1-\frac{2Mr^{2}}{r^{3}+\ell\left(1-\ell r\right)}\,.$ (13)
To the best of our knowledge, this Hayward-like RBH has never been proposed before. Just as the Hayward BH, it is evident that this space-time possesses a dS core, with an effective cosmological constant given by $\Lambda=6M/\ell$. We explicitly verify the finiteness of the curvature invariants, which take the following very simple forms:
$R$ $=$ $g^{\mu\nu}R_{\mu\nu}=\frac{24M}{\ell}\,,\quad R_{\mu\nu}R^{\mu\nu}=\frac{144M^{2}}{\ell^{2}}\,,$
${\cal K}$ $=$ $R_{\mu\nu\sigma\rho}R^{\mu\nu\sigma\rho}=\frac{96M^{2}}{\ell^{2}}$ (14)
We later explicitly verify that the intensity of Hawking radiation of the Hayward-like RBH increases with increasing regularizing parameter, therefore tightening constraints on $f_{\text{pbh}}$, and further closing the asteroid mass window. This behavior is similar to that of Kerr BHs, which display an enhancement of the primary photon emission spectrum with increasing spin parameter Arbey et al. (2020a, b, 2021b). A similar behaviour has also been observed in the context of the Kazakov-Solodukhin BH Kazakov and Solodukhin (1994) in Ref. Bronnikov and Konoplya (2023), consistently with the expectation laid out in footnote 3 (note, in fact, that the shadow of the Kazakov-Solodukhin BH increases with increasing regularizing parameter, see e.g. Fig. 1 of Ref. Kocherlakota et al. (2021), Fig. 18 of Ref. Akiyama et al. (2022), and Fig. 9 of Ref. Vagnozzi et al. (2023)). Finally, we have checked that the allowed limit for the regularizing parameter in terms of the horizon radius is $0\leq\ell\leq r_{H}$.
@@ III.1 Greybody factors
In general, numerical integration methods are required to compute GBFs for general space-times, and this holds for our tr-symmetric RBHs as well. In our work, we make use of the so-called shooting method (see Appendix A for further details), which has already been successfully applied to these types of calculations in several earlier works Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2023); Calzà and Rosa (2023a, b).
@@
In order to further simplify the problem, we work in units of horizon radius and therefore set $r_{H}=1$, so that $r=x+1$. In these units, the metric functions of the four RBHs under consideration are given by the following:
$\begin{split}&f_{\text{B}}(x)=1-\frac{(1+\ell^{2})^{3/2}(x+1)^{2}}{\left(\ell^{2}+(x+1)^{2}\right)^{\frac{3}{2}}}\,,\\ &f_{\text{H}}(x)=1-\frac{(x+1)^{2}}{(1-\ell^{2})\left((x+1)^{3}-\frac{\ell^{2}}{\ell^{2}-1}\right)}\,,\\ &f_{\text{CGSV}}(x)=1-\frac{e^{\ell-\frac{\ell}{x+1}}}{x+1}\,,\\ &f_{\text{H-l}}(x)=1-\frac{(1+\ell-\ell^{2})(x+1)^{2}}{\left((x+1)^{3}+\ell-\ell^{2}(x+1)\right)}\,,\end{split}$ (24)
for the Bardeen, Hayward, CGSV, and Hayward-like space-times respectively.
@@
where $\Gamma$ is the $\Gamma$ function. Using the method discussed above, we compute the GBFs for perturbations of different spin on the backgrounds of the four regular metrics discussed earlier, for different values of the regularizing parameter $\ell$. In the specific case $s=1$, we have checked that calculating the GBFs up to $l=4$ is sufficient for our purposes. The GBFs we calculate are then used to characterize the Hawking evaporation spectra.
@@ III.2 Evaporation spectra
Figure 2: Primary photon spectra resulting from the evaporation of Bardeen black holes of mass $10^{13}\,{\text{kg}}$ for different values of the regularizing parameter $\ell$ (normalized by the horizon radius $r_{H}$): $\ell/r_{H}=0.15$ (red dotted curve), $0.3$ (green dashed curve), and $0.45$ (magenta dash-dotted curve). The blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole.
@@
Figure 5: As in Fig. 2, but for Hayward-like black holes, with values of the regularizing parameter $\ell/r_{H}=0.15$ (red dotted curve), $0.3$ (green dashed curve), $0.45$ (magenta dash-dotted curve), $0.6$ (black dash-dotdotted curve), and $0.75$ (brown dash-dotdotdotted curve), whereas the blue solid curve corresponds once more to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole.
@@
Contrarily to what we observe for these three PRBHs, we see that the intensity of the spectra for Hayward-like PRBHs substantially increase with increasing $\ell$ for sufficiently large energies $E\gtrsim{\cal O}({\text{MeV}})$, whereas for lower energies they very slightly decrease. This behaviour agrees qualitatively with the expectations which motivated us to construct the Hayward-like RBH in first place, and with the temperature evolution shown in Fig. 1. In addition, we observe that an increase in the regularizing parameter leads to a much richer structure in the spectra: the observed multi-peak structure is the result of the contribution of different $l$ modes in the emission being disentangled more clearly. A similar behaviour is actually observed in the evaporation spectra of Kerr BHs which, by virtue of their rotation, lead to a more complex spectrum where the contribution of high-$l$ modes, while subdominant with respect to the $l=1$ mode, emerges more clearly (although we have explicitly checked that adding $l>4$ modes has virtually no effect on our results). In the case of Hayward-like RBHs, we remain agnostic as to the physical reason why such a structure is observed, and leave a more detailed study to follow-up work.
@@ III.3 Evaporation constraints
The spectra calculated in Sec. III.2 are then used to set evaporation constraints on $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, the fraction of DM in the form of PBHs, where $\Omega_{\text{pbh}}$ and $\Omega_{\text{dm}}$ are the PBH and DM density parameters respectively. Specifically, the computed spectra are used to obtain predictions for the flux of photons resulting from Hawking evaporation, which are then directly compared against measurements of the extragalactic photon background across a wide range of energies (see e.g. Ref. Auffinger (2023) for a recent review). Evaporation constraints are the dominant ones in the $10^{10}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{15}\,{\text{kg}}$ mass range: the lower limit of the range is set by the requirement that PBHs have not yet evaporated at the time of recombination, whereas the upper limit is defined by measurements of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, given that the intensity of the Hawking radiation flux is inversely proportional to the mass of the evaporating BH. In what follows, we will direct our attention exclusively to PBHs for which $M_{\text{pbh}}\gtrsim 10^{12}\,{\text{kg}}$: these have yet to fully evaporate today and, having formed deep during the radiation domination era, are therefore excellent non-baryonic DM candidates.
@@
We work under the commonly adopted assumption that PBHs are isotropically distributed on sufficiently large scales. Therefore, the flux resulting from their evaporation and reaching us today is given by the redshifted sum of the contributions from all evaporating PBHs in our Universe, and can be used to constrain the average extragalactic distribution of DM in the form of PBHs. Furthermore, we work within the (also commonly adopted) approximation of monochromatic mass distributions (which can be expected if the formation mechanism arises from an amplification of the power spectrum at a very specific scale), although the effect of extended mass distributions is the subject of active research Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b). Finally, as discussed earlier, we only consider the primary photon contribution, as the secondary component resulting from the decay into photons of other unstable particles is verified a posteriori to be negligible given the mass range of interest. While all these are clearly approximations, albeit widely adopted ones, we are confident that they are appropriate given the aim of our work. Our main goal is in fact to examine how the limits on $f_{\text{pbh}}$ change when moving from the Schwarzschild PBH framework to that of the regular metrics presented in Sec. II, potentially opening or closing the asteroid mass window. It is more than reasonable to expect that the shift in constraints relative to the Schwarzschild case, $\delta f_{\text{pbh}}$, is only weakly affected by the above approximations. In other words we expect these approximations to have similar impacts on the constraints on $f_{\text{pbh}}$ relative to Schwarzschild BHs for the different RBHs discussed in Sec. II, therefore leading to negligible effects on the relative shift $\delta f_{\text{pbh}}$, which is the quantity we are ultimately interested in. At any rate, the adopted approximations also allow for a more direct comparison to several previous works, and we therefore consider them appropriate for our pilot study, while stressing that their impact should definitely be explored in future follow-up works. Finally, note that we are tacitly assuming that PBHs cluster in the galactic halo in the same way as other forms of DM (unless they are extremely large, which is not the case for the mass range of interest).
@@
In what follows, we therefore assume that PBHs all have the same initial mass $M_{\text{pbh}}$. Following Ref. Carr et al. (2010) we approximate the number of emitted photons in the logarithmic energy bin $\Delta E_{\gamma}\simeq E_{\gamma}$ as being given by $\dot{N}_{\gamma}(E_{\gamma})\simeq E_{\gamma}(d\dot{N}_{\gamma}/dE_{\gamma})$. The emission rate of photons from Hawking evaporation per volume at a cosmological time $t$ is then given by Carr et al. (2010): [^5] In the original paper $M_{\text{pbh}}$ is assumed to be a function of time, due to the evaporation process. However, for the PBH masses considered in the present work ($M>10^{12}$kg) we can safely assume $M$ to be roughly constant during the evaporation process Carr and Kuhnel (2022).
@@
We assume a spatially flat $\Lambda$CDM cosmological model in specifying the expansion rate entering into Eq. (31), with the same cosmological parameters as in Ref. Carr et al. (2010). This allows us to robustly cross-check our Schwarzschild constraints on $f_{\text{pbh}}$ against those reported in the seminal Ref. Carr et al. (2010), although we stress that our constraints are very stable against reasonable changes in the values of the cosmological parameters. Once the cosmological model is fixed, all the relevant quantities in Eq. (31) are known except for the present-day PBH number density, $n_{\text{pbh}}(t_{0})$, which can be constrained from EGRB observations and is ultimately related to $f_{\text{pbh}}$. More specifically, for any given value of the PBH mass $M_{\text{pbh}}$, through Eqs. (29,31,32) we can compute the unnormalized photon flux $I(E_{\gamma 0})/n_{\text{pbh}}(t_{0})$, and adjust the normalization $n_{\text{pbh}}(t_{0})$ by comparing against EGRB observations (as we will explain shortly). This procedures gives us an upper limit on $n_{\text{pbh}}(t_{0})$, which can be translated into an upper limit on $f_{\text{pbh}}$ as follows:
@@
Figure 6: Photon fluxes resulting from the evaporation of primordial Bardeen black holes with regularizing parameter $\ell=0.3r_{H}$, and masses of $M_{\text{pbh}}=10^{12}\,{\text{kg}}$ (brown curve), $2.3\times 10^{12}\,{\text{kg}}$ (purple curve), $5\times 10^{12}\,{\text{kg}}$ (red curve), $10^{13}\,{\text{kg}}$ (green curve), $3\times 10^{13}\,{\text{kg}}$ (green curve), and $5\times 10^{13}\,{\text{kg}}$ (blue curve). The triangles indicate $1\sigma$ upper limits on the extragalactic $\gamma$-ray background flux as measured by HEAO-1 (maroon), COMPTEL (turquoise), and EGRET (teal). For each value of $M_{\text{pbh}}$, the PBH fraction $f_{\text{pbh}}$ has been set to its upper limit, determined as soon as the first datapoint is overshot (the latter is different for different values of $M_{\text{pbh}}$).

### v3

@@ II Regular black holes
A final caveat is in order before discussing the RBH metrics we consider. The latter are all regular in the sense of having finite curvature invariants $R\equiv g^{\mu\nu}R_{\mu\nu}$, $R_{\mu\nu}R^{\mu\nu}$, and ${\cal K}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. However, a more stringent criterion for regularity is that of geodesic completeness, which does not necessarily imply finiteness of curvature invariants and viceversa. We recall that a space-time is geodesically complete if any geodesic thereon can be extended to arbitrary values of their affine parameter. In other words, all geodesics extend (or can be extended) for all times. A number of “popular” RBHs have indeed been shown to have finite curvature invariants but to be geodesically incomplete Zhou and Modesto (2023). This includes the well-known Hayward RBH, which is among the ones we shall consider here. Nevertheless, given the significant interest in this metric, the fact that it is widely taken as prototype for RBHs, and our phenomenological goal of going beyond Schwarzschild PBHs, we will take this space-time into consideration, while cautioning the reader about the above issues, and therefore that the Hayward metric (but more generally all the metrics we will study) should be considered nothing more than phenomenological toy models at this stage. Note, in addition, that the stability of RBHs featuring inner horizons is currently a matter of debate in the literature Carballo-Rubio et al. (2021, 2022); Bonanno et al. (2023); Bonanno and Saueressig (2022); Carballo-Rubio et al. (2023).
@@ II.1 Bardeen black hole
where, in terms of the BH mass $M$, the regularizing parameter satisfies $\ell\leq\sqrt{16/27}M\sim 0.77\,M$ in order for the metric to describe a BH and not a horizonless object. Note that the Schwarzschild metric function is recovered in the $\ell\to 0$ limit. A perhaps physically more motivated choice is to express quantities in units of the horizon radius $r_{H}$, defined as the largest root of the equation $f(r_{H})=0$, in which case the regularizing parameter is subject to the constraint $\ell\lesssim 0.70\,r_{H}$ (which is the point at which $T=0$ in Fig. 1). In order to obtain this limit we have computed the solution to $f(r_{H})=0$ fixing $M=1$, in order to extrapolate $r_{H}(\ell)$, before analyzing for which real values of the parameter $n$ the equation $\ell=nr_{H}(\ell)$ admits solutions. We note that the peculiar factor of $\sqrt{16/27}$ in the extremality condition appears when one demands that the cubic equation determining the horizon(s) location(s) of the Bardeen RBH, $f_{B}(r)=0$, admits one real non-zero root: with some algebraic manipulation, one sees that this condition amounts to the requirement that $\sqrt{27\ell^{8}M^{4}-16\ell^{6}M^{6}}$ is real, from which the requirement $\ell\leq\sqrt{16/27}M$ follows.
@@ II.2 Hayward black hole
If expressed in terms of BH mass $M$, the regularizing parameter for the Hayward BH is subject to the same limit as that of the Bardeen BH, i.e. $\ell\leq\sqrt{16/27}\,,M$. On the other hand, if expressed in terms of the more physically motivated horizon radius, the limit is instead $\ell\lesssim 0.57\,r_{H}$ (again this is the point where $T=0$ in Fig. 1). We note that the Schwarzschild metric function is recovered in the $\ell\to 0$ limit. The factor of $\sqrt{16/27}$ in the extremality condition originates from considerations similar to those we made for the Bardeen RBH in Sec. II.1: with some algebraic manipulation, one sees that this condition amounts to the requirement that $\sqrt{27\ell^{4}M^{2}-16\ell^{2}M^{4}}$ is real, from which the requirement $\ell\leq\sqrt{16/27}M$ follows.
@@ III.1 Greybody factors
In general, numerical integration methods are required to compute GBFs for general space-times, and this holds for our tr-symmetric RBHs as well. In our work, we make use of the so-called shooting method (see Appendix A for further details), which has already been successfully applied to these types of calculations in several earlier works Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2024c); Calzà and Rosa (2024, 2023).
@@
In order to further simplify the problem, we work in units of horizon radius and therefore set $r_{H}=1$, so that $r=x+1$. In these units, the metric functions of the three RBHs under consideration are given by the following:
$\begin{split}&f_{\text{B}}(x)=1-\frac{(1+\ell^{2})^{3/2}(x+1)^{2}}{\left(\ell^{2}+(x+1)^{2}\right)^{\frac{3}{2}}}\,,\\ &f_{\text{H}}(x)=1-\frac{(x+1)^{2}}{(1-\ell^{2})\left((x+1)^{3}-\frac{\ell^{2}}{\ell^{2}-1}\right)}\,,\\ &f_{\text{CGSV}}(x)=1-\frac{e^{\ell-\frac{\ell}{x+1}}}{x+1}\,,\\ \end{split}$ (22)
for the Bardeen, Hayward, and CGSV space-times respectively.
@@
where $\Gamma$ is the $\Gamma$ function. Using the method discussed above, we compute the GBFs for perturbations of different spin on the backgrounds of the three regular metrics discussed earlier, for different values of the regularizing parameter $\ell$. In the specific case $s=1$, we have checked that calculating the GBFs up to $l=4$ is sufficient for our purposes. The GBFs we calculate are then used to characterize the Hawking evaporation spectra.
Figure 2: Greybody factors $\Gamma_{l=1}^{s=1}$ as a function of $\omega/M$ for Schwarzschild BHs (black curve), as well as the Bardeen (blue solid curve), Hayward (red dashed curve), and Culetu-Ghosh-Simpson-Visser (green dotted curve) regular space-times. For illustrative purposes we only plot $\Gamma_{l=1}^{s=1}$, since we are interested in photons ($s=1$) and the dominant emission mode is the $l=1$ one. We have fixed the regularizing parameter to $\ell=0.3r_{H}$ for all three regular space-times. We see that in all three cases the GBFs are consistently (slightly) higher than their Schwarzschild counterparts. The features shown in this plot do not change sensibly for higher values of $l$ and other values of $\ell$.
In Fig. 2 we show the $\Gamma_{l=1}^{s=1}$ GBFs for the Schwarzschild BH and the three regular space-times we study. For illustrative purposes we have focused on the $\Gamma_{l=1}^{s=1}$ GBFs, since we are interested in photons ($s=1$) and the dominant emission mode is the $l=1$ one. We also note that, since we are considering spherically symmetric space-times, the $(2l+1)$ different $m$ modes are degenerate. We have fixed the regularizing parameter to $\ell=0.3r_{H}$ for all three regular space-times, also for illustrative purposes. We see that, for all three regular space-times, the GBFs are slightly higher than their Schwarzschild counterparts (by $\lesssim 20\%$ at most), and asymptote to the latter for both $\omega/M\lesssim 0.3$ and $\omega/M\gtrsim 0.7$. While for definiteness we have focused on these specific values of $s$, $l$, and $\ell$, we have explicitly checked that the features described above are present for other values of the parameters in question as well.
@@ III.2 Evaporation spectra
Figure 3: Primary photon spectra resulting from the evaporation of Bardeen black holes of mass $10^{16}\,{\text{g}}$ for different values of the regularizing parameter $\ell$ (normalized by the horizon radius $r_{H}$): $\ell/r_{H}=0.15$ (red dotted curve), $0.3$ (green dashed curve), and $0.45$ (magenta dash-dotted curve). The blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole.
@@
For the Bardeen, Hayward, and CGSV PRBHs we observe that an increase in the regularizing parameter $\ell$ leads to a decrease in the intensity of the spectra at all energies. This is naïvely what one could expect by inspecting the temperature evolution shown in Fig. 1, since for all three these PRBHs the temperatures decrease relative to their Schwarzschild counterparts. However, as Eq. (27) shows, the temperature is not the only quantity at play, since the GBFs enter as well. As we have seen in Fig. 2, the GBFs for all three PRBHs are larger than their Schwarzschild counterparts, which would seem to counteract the effect of a lower temperature. A posteriori, however, the dominant effect turns out to be that of a lower temperature. We could actually have expected as much, since the resulting spectra given by Eq. (27) are linear in the GBFs, but depend exponentially on the temperature. Therefore, one might expect that a decrease in the temperature will have a much more dramatic effect than a comparable increase in the GBFs. This is precisely what we observe in Figs. 34, and 5.
Aside from the amplitude of the spectrum, for the Bardeen and Hayward PRBHs we note that the position of the peak in the spectrum is only mildly affected by the regularizing parameter, an increase in which pushes the peak towards slightly lower energies. On the other hand, for the CGSV BH an increase in the regularizing parameter pushes the peak towards slightly higher energies. We do not exclude that this different behaviour may be related to the type of core being considered, and we defer a more detailed investigation of this point to future work. At any rate, we expect that the behaviour of the spectra discussed above should lead to constraints on $f_{\text{pbh}}$ which are loosened for these classes of PRBHs.
@@ III.3 Evaporation constraints
The spectra calculated in Sec. III.2 are then used to set evaporation constraints on $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, the fraction of DM in the form of PBHs, where $\Omega_{\text{pbh}}$ and $\Omega_{\text{dm}}$ are the PBH and DM density parameters respectively. Specifically, the computed spectra are used to obtain predictions for the flux of photons resulting from Hawking evaporation, which are then directly compared against measurements of the extragalactic photon background across a wide range of energies (see e.g. Ref. Auffinger (2023) for a recent review). Evaporation constraints are the dominant ones in the $10^{13}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{18}\,{\text{g}}$ mass range: the lower limit of the range is set by the requirement that PBHs have not yet evaporated at the time of recombination, whereas the upper limit is defined by measurements of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, given that the intensity of the Hawking radiation flux is inversely proportional to the mass of the evaporating BH. In what follows, we will direct our attention exclusively to PBHs for which $M_{\text{pbh}}\gtrsim 10^{15}\,{\text{g}}$: these have yet to fully evaporate today and, having formed deep during the radiation domination era, are therefore excellent non-baryonic DM candidates. There is another important reason for focusing on this mass range. As shown in Appendix B, PBHs (either Schwarzschild or the three PRBHs we consider) within this range have lifetimes much longer than the age of the Universe, are far from having fully evaporated at the present time, and have only lost a negligible fraction of their mass from formation until now. When using the symbol $M_{\text{pbh}}$, we are therefore safe in denoting the values of the PBH mass both at formation and now.
@@
We work under the commonly adopted assumption that PBHs are isotropically distributed on sufficiently large scales. Therefore, the flux resulting from their evaporation and reaching us today is given by the redshifted sum of the contributions from all evaporating PBHs in our Universe, and can be used to constrain the average extragalactic distribution of DM in the form of PBHs. Furthermore, we work within the (also commonly adopted) approximation of monochromatic mass distributions (which can be expected if the formation mechanism arises from an amplification of the power spectrum at a very specific scale), although the effect of extended mass distributions is the subject of active research Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b). Finally, as discussed earlier, we only consider the primary photon contribution, as the secondary component resulting from the decay into photons of other unstable particles is verified a posteriori to be negligible given the mass range of interest. While all these are clearly approximations, albeit widely adopted ones, we are confident that they are appropriate given the aim of our work. Our main goal is in fact to examine how the limits on $f_{\text{pbh}}$ change when moving from the Schwarzschild PBH framework to that of the regular metrics presented in Sec. II, altering the asteroid mass window. It is more than reasonable to expect that the shift in constraints relative to the Schwarzschild case, $\delta f_{\text{pbh}}$, is only weakly affected by the above approximations. In other words we expect these approximations to have similar impacts on the constraints on $f_{\text{pbh}}$ relative to Schwarzschild BHs for the different RBHs discussed in Sec. II, therefore leading to negligible effects on the relative shift $\delta f_{\text{pbh}}$, which is the quantity we are ultimately interested in. At any rate, the adopted approximations also allow for a more direct comparison to several previous works, and we therefore consider them appropriate for our pilot study, while stressing that their impact should definitely be explored in future follow-up works. Finally, note that we are tacitly assuming that PBHs cluster in the galactic halo in the same way as other forms of DM (unless they are extremely large, which is not the case for the mass range of interest).
@@
In what follows, we therefore assume that PBHs all have the same initial mass $M_{\text{pbh}}$. Following Ref. Carr et al. (2010) we approximate the number of emitted photons in the logarithmic energy bin $\Delta E_{\gamma}\simeq E_{\gamma}$ as being given by $\dot{N}_{\gamma}(E_{\gamma})\simeq E_{\gamma}(d\dot{N}_{\gamma}/dE_{\gamma})$. The emission rate of photons from Hawking evaporation per volume at a cosmological time $t$ is then given by Carr et al. (2010): [^5] In the original paper $M_{\text{pbh}}$ is assumed to be a function of time, due to the evaporation process. However, for the PBH masses considered in the present work ($M>10^{15}\,{\text{g}}$) we can safely assume $M$ to be roughly constant during the evaporation process Carr and Kuhnel (2022), see Appendix B for more detailed discussions.
@@
We assume a spatially flat $\Lambda$CDM cosmological model in specifying the expansion rate entering into Eq. (29), with the same cosmological parameters as in Ref. Carr et al. (2010). This allows us to robustly cross-check our Schwarzschild constraints on $f_{\text{pbh}}$ against those reported in the seminal Ref. Carr et al. (2010), although we stress that our constraints are very stable against reasonable changes in the values of the cosmological parameters. Once the cosmological model is fixed, all the relevant quantities in Eq. (29) are known except for the present-day PBH number density, $n_{\text{pbh}}(t_{0})$, which can be constrained from EGRB observations and is ultimately related to $f_{\text{pbh}}$. More specifically, for any given value of the PBH mass $M_{\text{pbh}}$, through Eqs. (27,29,30) we can compute the unnormalized photon flux $I(E_{\gamma 0})/n_{\text{pbh}}(t_{0})$, and adjust the normalization $n_{\text{pbh}}(t_{0})$ by comparing against EGRB observations (as we will explain shortly). This procedures gives us an upper limit on $n_{\text{pbh}}(t_{0})$, which can be translated into an upper limit on $f_{\text{pbh}}$ as follows:
@@
Figure 6: Photon fluxes resulting from the evaporation of primordial Bardeen black holes with regularizing parameter $\ell=0.3r_{H}$, and masses of $M_{\text{pbh}}=10^{15}\,{\text{g}}$ (brown curve), $2.3\times 10^{15}\,{\text{g}}$ (purple curve), $5\times 10^{15}\,{\text{g}}$ (red curve), $10^{16}\,{\text{g}}$ (green curve), $3\times 10^{16}\,{\text{g}}$ (green curve), and $5\times 10^{16}\,{\text{g}}$ (blue curve). The triangles indicate $1\sigma$ upper limits on the extragalactic $\gamma$-ray background flux as measured by HEAO-1 (maroon), COMPTEL (turquoise), and EGRET (teal). For each value of $M_{\text{pbh}}$, the PBH fraction $f_{\text{pbh}}$ has been set to its upper limit, determined as soon as the first datapoint is overshot (the latter is different for different values of $M_{\text{pbh}}$).